RHIC Running: NSRL started running experiments this morning. No issues.
In RHIC, the Beam intensity is up. Polarization down. Did a scan in the AGS of intensity vs polarization, and it was not possible to get more than 50% at these intensities. 53/47 at injection, down to 51/38 at store. Yesterday's beams were 1.48/1.43e11 with 11.5kHz PHENIX ZDCs. This morning we ran 1.61/1.56e11 at 12.5kHz, decaying within 15 minutes to 11kHz. Polarization is 54/53% at injection and 53/40% at store. The last two stores appear to show the polarization in Yellow is no longer dropping throughout the store. There is no known change in configuration.
STAR reports that the electronics on the front of the TPC is failing, due probably to high radiation levels from these high beam intensities.
Today, after the morning fill we will swap the orbit of the Jet from Blue to Yellow. Yun will work on the nonlinear chromaticity correction. This will take the beam for ~2 hours. The Blue beam synch was lost in the morning store, so there will be a short access to allow this to be reset. JET requires a half hour access for service. Before the access, the JET requires a 4 hour warm up, then access, and then a 4 hour cool down. Currently the JET is taking off-axis data, and they do not want to do the servicing until after this fill, and probably not before tomorrow.
BLIP is going to 24/7 operation. Peter Ingrassia will ensure there is operator/CAS coverage..
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |